# Regularity and temperature of stationary black hole event horizons

### v1

@@ 5 Conclusions
Ultimately, by using the smoothness of the horizon, we have been able to prove the existence of cross-sections, but the logical derivation was reversed. Finally, we obtained a theorem that describes in detail the geometry of the black hole horizon. We showed the possibility of identifying the surface gravity itself (not just its sign), something that allows its identification with the temperature, and we also identified the angular velocity proving the existence of an axisymmetry. We showed that these quantities are well defined even without imposing the Einstein’s equation, just on the basis of the horizon geometry.
It is likely that our methods could be applied, suitably adapted, to the study of higher dimensional black holes and certainly to studies in the analytic setting [26, 25]. We leave these research directions for future work.
Summarizing, stressing the physically relevant conclusions, we showed the constancy of surface gravity and so proved the zeroth’s law of black hole thermodynamics without assumptions on (a) smoothness of the horizon, (b) non-degeneracy, (c) existence of cross-section, (d) (electro-)vacuum conditions and under (e) fairly weak causality condition conditions (strong causality). Only stationarity was substantially used but this is necessary if one wants to derive a constant temperature, as it expresses the fact that the black hole has reached thermal equilibrium.

### v2

@@ 8 Conclusions
Ultimately, by using the smoothness of the horizon, we have been able to prove the existence of cross-sections, but the logical derivation was reversed. Finally, we obtained a theorem that describes in detail the geometry of the black hole horizon. We showed the possibility of identifying the surface gravity itself (not just its sign), something that allows its identification with the temperature, and we also identified the angular velocity proving the existence of a horizon axisymmetry. We showed that these quantities are well defined even without imposing the Einstein’s equation, just on the basis of the horizon geometry.
It is likely that our methods could be applied, suitably adapted, to the study of higher dimensional black holes and certainly to studies in the analytic setting hollands07 ; hollands09 . We leave these research directions for future work.
Summarizing, stressing the physically relevant conclusions, we showed the constancy of surface gravity and so proved the zeroth’s law of black hole thermodynamics without (a) smoothness assumptions on the horizon, (b) non-degeneracy conditions, (c) hypothesis on the existence of cross-sections, (d) (electro-)vacuum conditions (e) horizon bifurcation assumptions, and under fairly weak causality conditions (strong causality). Only stationarity was substantially used but this is necessary if one wants to derive a constant temperature, as it expresses the fact that the black hole has reached thermal equilibrium.
@@ Acknowledgments
Declarations
Data availability statement
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
Conflicts of Interest
The authors have no relevant financial or non-financial interests to disclose.